Calculate the binding energy per nucleon in MeV for the iron isotope with an atomic
number of 26 and a mass number of 56,
56Fe26 which has an atomic mass of 55.9349 amu.
1
Expert's answer
2019-07-15T07:04:37-0400

•     Binding Energy: The amount of energy released when a nucleus is formed from its components.•        Mass defect ΔM = Z*Mp + N*Mn-MAZ= Atomic no., N= No. of neutrons Mp = Mass of proton, Mn = Mass of neutron,MA= Atomic Mass


Binding Energy = ΔMc2 = ΔM * 931.5 MeV


Binding Energy Per Nucleon = ΔMc2 = (ΔM * 931.5 MeV)/ (no. of protons + no. of neutrons)


Here ΔM can be calculated from above equation as:


Z=26, N=30, Mp=1.0078 u, MN= 1.0086 u.
ΔM = 26(1.0078) +30(1.0086)-55.9349ΔM = 0.5259 u


Binding Energy= ΔMc2 = 0.5259 * 931.5 MeV = 489.87 MeV


Binding Energy Per Nucleon = (489.87)/56 = 8.74 MeV